# Surrogate-Assisted Targeted Learning for Delayed Outcomes under Administrative Censoring

### 11 Mar 2026 v1

@@ 7 Design-Calibrated Illustration: Washington State EPT Trial
The EPT trial (Golden et al., 2015) assessed whether a public health programme promoting free patient-delivered partner therapy could reduce chlamydia positivity among women aged 14–25 in Washington State. The four-wave stepped-wedge design randomized $J=23$ local health jurisdictions (LHJs) across $T=5$ time periods (one baseline, four post-baseline steps of six to eight months each): six LHJs crossed over at each of waves 1–3 and five at wave 4, with approximately 162 chlamydia tests per LHJ per period at sentinel clinical sites. The delayed primary outcome is 12-month chlamydia test positivity. This is administratively censored for late-crossing LHJs: wave-4 LHJs ($t_{j}=4$) have under 12 months of post-crossover follow-up, yielding an administrative censoring rate of 86% in that wave; wave-3 LHJs are partially censored (40%) from period 3 onward; overall 33.7% of records have missing 12-month outcomes. EPT uptake — whether the index patient accepted patient-delivered partner therapy within 30 days of STI diagnosis — is complete for all subjects regardless of wave and satisfies the three surrogate criteria of Remark 2: causal proximity to the intervention, temporal availability before the 12-month outcome, and biological stationarity.
@@ 8 Discussion
We proposed a Surrogate-Assisted TMLE linking the longitudinal G-computation formula (1) to the cluster-level EIC of Section 3, providing an asymptotically doubly robust, non-parametric solution to the administrative censoring of late-crossing clusters in SW-CRTs. Standard IPCW fails in this setting because the censoring propensity $g_{\Delta}$ can become arbitrarily small for late-crossing clusters (Remark 3), producing highly unstable weights. The Surrogate-Assisted TMLE bypasses this instability by bridging missing outcomes through $S$, which is observed for all individuals regardless of the censoring horizon; the simulation study of Section 6 confirms near-zero bias and markedly lower variance than IPCW across all scenarios evaluated.
@@
The observed CV-TMLE coverage of 0.87–0.91 highlights a finite-sample regime in which the asymptotic approximation is directionally informative but not yet fully calibrated. Our finite-sample experiments should be read as showing qualitative agreement with the asymptotic theory rather than exact small-sample calibration; in the regimes studied here, the remaining undercoverage appears driven by nuisance-estimation error that is not yet negligible at the available cluster counts. Table 5 documents that the empirical variance exceeds the mean sandwich by a factor of $\approx 1.6$ at all $J$, qualitatively consistent with the boundary case $\alpha+\beta\approx 1/2$ of Corollary S2 (Supplement, Section S2.7). The $t_{J-1}$ interval is a practical safeguard, but not a full cure; practical remedies include a richer Super Learner library, a cluster bootstrap, or the variance inflation correction described in Supplement, Section S2.7.
An important methodological implication of Proposition 2 is that the nested G-computation formula (1) cannot be targeted by a straightforward DML one-step without an additional rate condition on the conditional surrogate density $f_{S}$. The SA-TMLE’s two-stage nested fluctuation eliminates this requirement by constructing $\hat{\bar{Q}}^{*}_{\mathrm{int}}$ as a propensity-weighted empirical mean of $\hat{\bar{Q}}^{*}_{Y}$, marginalizing over the observed surrogate distribution without estimating $f_{S}$.
@@
The validity of the estimator hinges on Assumption 3 (Surrogate-Mediated MAR), whose plausibility is discussed in Remark 2. Assumption 3 is partially testable on the fully-observed sub-cohort via the augmented regression:
$\mathbb{E}[Y\mid S,A,W,t,\Delta=1]\;=\;\mathbb{E}[Y\mid S,A,W,t,\hat{r}(A,W,t),\Delta=1],$ (51)
where $\hat{r}(A,W,t)=\hat{P}(\Delta=1\mid A,W,t)/\hat{P}(\Delta=1\mid S,A,W,t)$; a zero coefficient on $\hat{r}$ is consistent with Assumption 3.
Sensitivity analysis under MAR violation.
When Assumption 3 fails, the bias of $\hat{\Psi}_{\mathrm{TMLE}}$ can be characterized via a tipping-point analysis. Let $\gamma(S,A,W,t)$ denote the magnitude of the MAR violation:
$\gamma(S,A,W,t)\;\equiv\;\mathbb{E}[Y\mid S,A,W,t,\Delta=0]\;-\;\mathbb{E}[Y\mid S,A,W,t,\Delta=1].$ (52)
This is the conditional mean difference in $Y$ between the unobserved ($\Delta=0$) and observed ($\Delta=1$) sub-cohorts, given the surrogate. Under Assumption 3, $\gamma\equiv 0$. Under a violation of magnitude $\gamma$, the leading-order bias of $\hat{\Psi}_{\mathrm{TMLE}}$ is:
$\mathrm{Bias}(\hat{\Psi}_{\mathrm{TMLE}})\;\approx\;\mathbb{E}\!\left[\frac{I(A=1)(1-g_{\Delta})}{g_{\Delta}}\,\gamma(S,1,W,t)\;-\;\frac{I(A=0)(1-g_{\Delta})}{g_{\Delta}}\,\gamma(S,0,W,t)\right],$ (53)
where the expectation is over the marginal distribution and $1-g_{\Delta}$ is the censoring probability. Equation (53) reveals two sensitivity drivers: the censoring rate $(1-g_{\Delta})/g_{\Delta}$ (large for late-crossing clusters) and the differential violation $\gamma(S,1,W,t)-\gamma(S,0,W,t)$ across arms. The tipping-point $\gamma^{*}$ is obtained by solving $\hat{\Psi}_{\mathrm{TMLE}}=\hat{\mathrm{Bias}}(\gamma^{*})$; values implausibly large given subject-matter knowledge indicate a robust conclusion. This extends Scharfstein et al. (1999) to the surrogate-mediated setting and is analogous to the $E$-value of VanderWeele and Ding (2017); we recommend reporting $\gamma^{*}$ routinely.
More broadly, the results here suggest that surrogate-assisted targeted learning can serve as a template for a wider class of nested semiparametric functionals in which a primary outcome is observed sparsely but an intermediate variable remains broadly available. Although the bridge formulation extends beyond stepped-wedge trials, the present theoretical and numerical development is calibrated to the stepped-wedge setting; extension to other longitudinal designs is a promising but separate task. The key lesson is not only that surrogate information can stabilize inference under administrative censoring, but that the geometry of the nested bridge functional itself may require estimator constructions beyond standard one-step debiasing. Extending these ideas to multistage treatments, informative surrogate measurement, and more general longitudinal bridge structures is a natural direction for future work.
8.1 Future Directions
Two directions remain: (i) extending the cluster-level EIC to time-to-event outcomes with continuous administrative censoring; and (ii) establishing non-asymptotic coverage bounds for the CV-TMLE when nuisance classes are unbounded, which requires bracketing entropy bounds beyond those currently available for Super Learner ensembles. The $t_{J-1}$ confidence interval (41) provides a practical safeguard, formally justified by Theorem S1.

### 15 Mar 2026 v2

@@ 7 Design-Calibrated Illustration: Washington State EPT Trial
The EPT trial (Golden et al., 2015) assessed whether a public health programme promoting free patient-delivered partner therapy could reduce chlamydia positivity among women aged 14–25 in Washington State. The four-wave stepped-wedge design randomized $J=23$ local health jurisdictions (LHJs) across $T=5$ time periods (one baseline, four post-baseline steps of six to eight months each): six LHJs crossed over at each of waves 1–3 and five at wave 4, with approximately 162 chlamydia tests per LHJ per period at sentinel clinical sites. The delayed primary outcome is 12-month chlamydia test positivity. This is administratively censored for late-crossing LHJs: wave-4 LHJs ($t_{j}=4$) have under 12 months of post-crossover follow-up, yielding an administrative censoring rate of 86% in that wave; wave-3 LHJs are partially censored (40%) from period 3 onward; overall 33.7% of records have missing 12-month outcomes. EPT uptake — whether the index patient accepted patient-delivered partner therapy within 30 days of STI diagnosis — is complete for all subjects regardless of wave and satisfies the three surrogate criteria of Assumption 3: causal proximity to the intervention, temporal availability before the 12-month outcome, and biological stationarity.
@@ 8 Discussion
We developed a Surrogate-Assisted TMLE for the nested bridge functional $\Psi(P_{0})=\mathbb{E}_{W,t}[\mathbb{E}_{S|A=1}[\bar{Q}_{Y}(S,1,W,t)]-\mathbb{E}_{S|A=0}[\bar{Q}_{Y}(S,0,W,t)]]$ under administrative censoring. The key theoretical contribution is the identification and resolution of $R_{SY}$ (Proposition 1): a second-order remainder with no doubly-robust complement that makes a standard DML one-step insufficient for this functional class. The two-stage nested fluctuation absorbs $R_{SY}$ without estimating the conditional surrogate density $f_{S}$, recovering $\sqrt{J}$-consistent asymptotic linearity under a single product-rate condition. Simulations confirm near-zero bias and markedly lower variance than IPCW across all censoring regimes evaluated.
@@
The sandwich interval for CV-TMLE undercovers at 0.87–0.89 across all cluster counts. Table 1 identifies the source: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but omits the variance of the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which contribute an additional $O(J^{-1})$ term that does not vanish faster than the leading sandwich term under the present nuisance specification. This is not a failure of the point estimator — bias is essentially zero at all $J$ — but a variance-calibration gap attributable to the two-stage targeting construction.
The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), resolves this gap by capturing the fluctuation-step variance the sandwich misses. Coverage is 0.956 at $J=30$ and 0.960–0.968 at $J\geq 50$ — near nominal across the practically important regime. At very small $J$ (10–20) the jackknife is mildly conservative (0.970–0.980); this is a well-understood property of leave-one-out procedures in small samples and represents a safe failure mode. We recommend the cluster jackknife as the default inference procedure for SA-TMLE, with the sandwich as a fast approximation when $J\geq 50$.
@@
Assumption 3 (Surrogate-Mediated MAR) is partially testable on the observed sub-cohort via the augmented regression $\mathbb{E}[Y\mid S,A,W,t,\hat{r}(A,W,t),\Delta=1]$, where $\hat{r}=\hat{P}(\Delta=1\mid A,W,t)/\hat{P}(\Delta=1\mid S,A,W,t)$; a zero coefficient on $\hat{r}$ is consistent with Assumption 3. When the assumption fails with violation $\gamma(S,A,W,t)\equiv\mathbb{E}[Y\mid S,A,W,t,\Delta=0]-\mathbb{E}[Y\mid S,A,W,t,\Delta=1]$, the leading-order bias is
$\mathrm{Bias}(\hat{\Psi}_{\mathrm{TMLE}})\;\approx\;\mathbb{E}\!\left[\tfrac{(1-g_{\Delta})}{g_{\Delta}}\bigl(I(A=1)\gamma(S,1,W,t)-I(A=0)\gamma(S,0,W,t)\bigr)\right].$ (42)
The tipping-point $\gamma^{*}$ solving $\hat{\Psi}_{\mathrm{TMLE}}=\widehat{\mathrm{Bias}}(\gamma^{*})$ extends Scharfstein et al. (1999) to the surrogate-mediated setting and is analogous to the $E$-value of VanderWeele and Ding (2017); we recommend reporting $\gamma^{*}$ routinely.
Scope and future work.
The surrogate-bridge framework applies wherever a delayed primary outcome is observed sparsely and an intermediate variable is broadly available; extension to other longitudinal designs beyond SW-CRTs is straightforward in principle. Two open directions are: (i) extending the cluster-level EIC to time-to-event outcomes with continuous administrative censoring; and (ii) establishing non-asymptotic coverage bounds for the CV-TMLE when nuisance classes are unbounded, requiring bracketing entropy bounds beyond those available for current Super Learner ensembles.